OpenStudy (thatcreepyboy):

Medal and fan. Find the height of a square pyramid that has a volume of 12 cubic feet and a base length of 3 feet. 1.5 feet 2 feet 3 feet 4 feet

OpenStudy (will.h):

Volume of a square pyramid is V= a^2 * h/3 We know that V= 12 and a= 3 so substitute 12 = 3^2 * h/3 12 = 9 * h/3 Divide both sides by 9 12/9 = h/3 1.33 = h/3 Multiply 3 by both sides 3 * 1.33 = h 4 = h

OpenStudy (thatcreepyboy):

Thank you.